When a test charge is placed at the centre of dipole then what will be the net force on the test charge? Please give good explanation.

Grade:12

1 Answers

Ash
5 years ago
Let two charges of equal magnitude Q but opposite sign be placed at 2a from each other.
Now a charge q0 is placed midway between them.
The force experienced by it due the negative charge is 
F=\frac{1}{4\pi \varepsilon}\frac{Qq_{0}}{a^{2}} towards the negative charge
The force experienced by it due the positive charge is 
F=\frac{1}{4\pi \varepsilon}\frac{Qq_{0}}{a^{2}} away from the postive charge
So,the net force experienced by it will be
F_{net}=2 \frac{1}{4\pi \varepsilon}\frac{Qq_{0}}{a^{2}} towards the negative charge.